OpenStudy (lilsis76):

arithmetic sequcences and series the fourth term is an arithmetic sequcence -6, and the 10th term is 5. find the common difference and the first term. how do i go about doing this? i tried so many numbers and the closest thing i got was 1/2 for the common difference :/

OpenStudy (sirm3d):

try this formula for an arithmetic sequence: \[\large a_n-a_k=d(n-k)\] use n=10, k=4, a10=5 and a4=-6 to solve the common difference

OpenStudy (anonymous):

terms are \[-6,-6+d,-6+2d,-6+3d,-6+4d,-6+5d,-6+6d=5\] solve for \(d\)

OpenStudy (sirm3d):

after solving d, use the same formula with n=10, a10=5, k=1 and the value of d to solve a1.

OpenStudy (lilsis76):

okay @sirm3d i got the a10=5 i get 2 and a4=-6 i got about -1 1/2

OpenStudy (anonymous):

hmm i get \[-6+6d=5\] \[6d=11\] \[d=\frac{11}{6}\]

OpenStudy (lilsis76):

wait, what is the formula i should be using? i have this one but i dont know how to use it: avn = av1 + (n - 1)d

OpenStudy (sirm3d):

you can also use that formula. take n=10 and av10=5 and you'll have two unknowns av1 and d use k=4 and av4=-6 and you'll get another equation in av1 and d.

OpenStudy (sirm3d):

\[\large n=10: av_{10} =av_1+(10-1)d\]\[\large n=4: av_{4}=av_1+(4-1)d\]

OpenStudy (lilsis76):

okay i get... 5= a + 9d and -6= a+3d

OpenStudy (lilsis76):

@sirm3d how can i figure out the rest?

OpenStudy (sirm3d):

multiply the second equation by (-1) then add it to the first equation. that should eliminate a and leave d to be solved.

OpenStudy (sirm3d):

\[\large \begin{matrix}5=a+9d \\ 6=-a-3d\end{matrix}\]

OpenStudy (lilsis76):

okay i get d=11/6

OpenStudy (lilsis76):

@sirm3d

OpenStudy (lilsis76):

is there anything else to do? i think there is but i dont know what

OpenStudy (lilsis76):

@Outkast3r09 is there anything i gotta do next ? and how?

OpenStudy (lilsis76):

@timo86m

OpenStudy (sirm3d):

now use d = 11/6 in either equation 1 or 2, then solve a \[\huge 5=a+9(\frac{11}{6})\]

OpenStudy (lilsis76):

i get -1 1/2

OpenStudy (sirm3d):

\[\large 5 - \frac{33}{2} = a\]

OpenStudy (sirm3d):